OpenStudy (countrygirl1431):

What two numbers add to make 36 with one of the numbers being exactly twice the other? smaller number = and larger number =

OpenStudy (jango_in_dtown):

Let the smaller number be x Then larger number =2x. By the given condition x+2x=36 so 3x=36 so x=12

OpenStudy (countrygirl1431):

so the larger number is 36?

OpenStudy (jango_in_dtown):

no. smaller =x=12. so larger =2x=2*12=???

OpenStudy (countrygirl1431):

24?

OpenStudy (jango_in_dtown):

correct

OpenStudy (countrygirl1431):

Fill in the blank to make the statement true. On Mr. Jones' farm there were exactly 6 cows for every pig. Mr. Jones had _______pigs. He had 114 cows.

OpenStudy (countrygirl1431):

would the answer be 19?

OpenStudy (jango_in_dtown):

right:)

OpenStudy (countrygirl1431):

thanks